A particle P is let fall vertically from a height to the ground . Another particle Q is projected vertically down 3s later with a velocity of 3m/s from the same height . If both particle reached the ground simultaneously , calculate 

(a)time by each particle to reach the ground

(d) distance travelled by each particle to the ground


1
Expert's answer
2021-02-17T10:41:08-0500

"\text{The formula for the magnitude of the displacement }\newline
\text{of the body thrown vertically down}:"

"h =V_0t+\frac{gt^2}{2}"

"\text{for particle P:}"

"V_0=0;h=\frac{gt^2}{2}"

"h =\frac{gt^2}{2}"

"\text{In 3 s, particle A:}"

"h_1 = 4.5g\text{ - distance covered in 3 }"

"V_1 =3g -\text{speed A through 3s}"

"\text{then:}"

"h = h_1+V_1t_f+\frac{gt^2_f}{2}\text{ where }"

"t_f =t-3;t- \text{total fall time of particle A}"

"h = 4.5g+3gt_f+\frac{gt^2_f}{2}"

"\text{for particle Q:}"

"V_0=3;h=3t_f+\frac{gt_f^2}{2}"

"\text{equate expression h for A and Q}"

"4.5g+3gt_f+\frac{gt^2_f}{2}= 3t_f+\frac{gt_f^2}{2}"

"t_f=\frac{4.5g}{3-3g}<0"

"t_f<0 \text{this is not a valid value}"

"\text{therefore the problem has no solution}"

Answer: For the given initial conditions, the problem has no solutions